My sister and I are having problems with the process of elimination.
We have to solve the story problem given:
The football team has 45 members. There are twice as many juniors as seniors. How many of each are on the team?

Answer by scott8148(6628)

"The football team has 45 members" ___ j + s = 45

"There are twice as many juniors as seniors" ___ j = 2s

subtracting the 2nd equation from the 1st (to eliminate j) ___ j + s - j = 45 - 2s

s = 45 - 2s ___ 3s = 45

solve for s, then substitute back to find j